Greenhab Report – January 12th

GreenHab Report

Mark Gee

12Jan2018

Environmental control:

Cooling with vent

Heating

Shade cloth on

Working Hour: 07:40PM
Inside temp at working hour: 19 C
Outside temp during working hours: 2 C
Inside temperature high: 31 C
Inside temperature low: 16 C
Inside humidity: 33 %RH

Inside humidity high: 49 %RH
Inside humidity low: 25 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: Harvested radish microgreens, lettuce, green beans, dill, and cilantro.

Daily water usage for crops: 22 gallons

Time(s) of watering for crops: 03:30PM

Research observations: Harvested microgreen experiment and for each treatment measured mass harvested, volume harvested, and size of plants.

Changes to research plants: Microgreen experiment is complete.

Aquaponics: Not in use.

Narrative: Today was harvest day! For the crops we harvested 580g of radish microgreens, a bowl of lettuce leaves, a fistfull of greenbeans, a handful of dill, and a sprinkle of cilantro. The other crops are growing well and should produce a bountiful harvest for crews to follow.

The microgreen experiment was harvested and has concluded successfully. It is too early for results, but keep an eye out for the upcoming report.

I have been making an effort to characterize the Greenhab and will be synthesizing my observations into a cohesive document. As a sneak peak, the figure tracking the temperature and humidity in the habitat throughout the day will be included in the Mission Summary as an image. 

Journalist Report – January 12th

Sol 12 – Harvest

Today was our final EVA. A small team consisting of Melanie Grande (crew engineer), Sam Albert (health and safety officer), and myself, departed the habitat around 11 am and embarked on a long and bumpy drive out to Skyline Rim. We did not arrive until well past noon. What initially seemed like a row of jagged teeth along the western horizon soon grew into a towering mesa of red stone that commanded our admiration. It was a humbling vista.

Our purpose was to collect and analyze samples of shale on behalf of our geologist, Cesare Guariniello. To do so, Sam and Melanie ascended an escarpment to the layers of stone strata. Though their height was hardly even half that of the monolithic ridge, Sam reported that he was able to see all the way to the habitat. But given the long drive back, we couldn’t stay for long. By 2:30 pm we had returned to the habitat with only a short break along the way.

Meanwhile, the rest of the crew has been busy measuring the yield of Mark’s microgreens experiment. Though most of the crop was reserved for his experiment, a portion was kept for tonight’s meal. The whole team gathered early this evening to partake in the harvest of fresh radish, lettuce, dill, cilantro, and green beans. It will be the first fresh food we’ve had since the mission began. For the graduate students on the team, the wait has been even longer!

The meal will be one to savor, for it will be our last dinner in this habitat. Mission control confirms that the refueling of our ascent vehicle using in-situ resources has completed and that the vehicle registers nominal on all system checkouts. Tomorrow at noon our crew will depart the habitat, travel to the ascent vehicle, and prepare for launch at dawn on Sunday morning. It will be the culmination of our grand adventure to the Red Planet. But until then, a frenzy of cleaning and preparation awaits!

Justin Mansell, MDRS Crew 186 Journalist

P.S. Photos attached. Photo of the day: 515300_4251200 12Jan2018 Skyline Rim.jpg.
(Sorry for the huge photo, but it really is incredible. The credit goes to Melanie Grande for taking it!)

EVA Report – January 12th

Purpose of EVA: Collect and analyze samples of shale obtained from the region below Skyline Rim. Collect samples of fossilized shells that were discovered on a previous EVA near White Moon.

Destination: Skyline Rim

UDM27 Coordinates: 515300E, 4251200N

Destination: White Moon

UDM27 Coordinates: 517100E, 4254600N

Participants: Justin Mansell (Journalist), Melanie Grande (Engineer), Sam Albert (Health and Safety Officer)

Narrative: This was the final EVA of the mission and the members of the team were tasked with obtaining samples of shale from the base of Skyline Rim and analyzing them in the field with a spectrometer. Our secondary objective was to collect a handful of fossilized shells from a deposit near White Moon. This deposit was discovered on a previous EVA but no samples were taken. We departed the airlock at 11:07 am and drove north to White Moon. I led in the Deimos rover followed behind by Sam and Melanie in Spirit. We made a brief stop near Gray Moon to search for the fossil bed but were unsuccessful in locating it. We remounted the rovers after several minutes and continued to Skyline Rim by way of Copernicus Highway 1574 and Sagan Street 1103. We arrived on site at approximately 12:30 pm.

Sam was able to make contact with the habitat from atop a small hill using a ham radio while Melanie and I searched for shale near the end of Sagan Street. Once we had positively identified several samples of shale the team collected equipment from the rovers and proceeded to the base of the cliff. We also made a brief test of the navigation radio to verify the direction to the habitat.

Sam made his way up an escarpment to collect chips of shale directly from the cliff wall using a rock hammer. Melanie also climbed an escarpment to take complimentary spectra using the spectrometer. I remained at the base of the escarpments to monitor Sam and Melanie’s safety. The team then returned to the rovers and departed Skyline Rim at 1:15 pm.

We made a second stop at White Moon on the return journey and relocated the shell fossils found on a previous EVA. We filled a sample bag with a handful of shells and returned directly to the habitat. We arrived at the habitat at 2:23 pm. The total time on EVA was 3 hours and 21 minutes.

Justin Mansell, MDRS Crew 186 Journalist

Operations Report – January 12th

Crew 186 Operations Report 12JAN2018

SOL: 12

Name of person filing report: M. Grande

Non-nominal systems: None

Notes on non-nominal systems: None.

Generator (hours run): 12h

Generator switched batteries, 7:30am

Generator turned off, charging battery at 10:30am

Generator turned on at 5:20pm

Solar— SOC

@ 7:30am : 40%

@ 10:30am : 100%

@ 5:20pm : 64%

Diesel: 50%

Propane: 26%

Ethanol Free Gasoline (5 Gallon containers for ATV): 3.3 Gallons

Water (trailer): 10 Gallons

Water (static): 316 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 129380.8 Gallons

Toilet tank emptied: No

ATVs Used: None

Oil Added? No

ATV Fuel Used: 00 Gallons

# Hours the ATVs were used today: 00:00 hours

Notes on ATVs: None.

Deimos rover used: Yes

Hours: 101.0

Beginning charge: 100%

Ending charge: 84%

Currently charging: Yes

Sojourner rover used: Assigned to director only.

Hours: 5.5

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Spirit rover used: Yes

Hours: 12.1

Beginning charge: 100%

Ending charge: 47%

Currently charging: Yes

Opportunity rover used: No

Hours: 5.5

Beginning charge: 100%

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 3.1

Beginning charge: 100%

Ending charge:

Currently charging: No, but was charging this morning during EVA

HabCar used and why, where? No

General notes and comments: Very fun EVA today because we had a nice long ride out to Skyline Rim, and I was just completely awed by the views. Sam and I each climbed up parts of the escarpment (or the more bougie term: “colluvial fan”) at the bottom of the ridge, trying to touch the layers left by Jurassic forces millions of years before us. Looking up to see the scale of one man compared to that ridge was incredible. Deimos broke 100 hours today– congratulations to our old chap! Young Spirit hung in there and is now at 12h; we’re slowly increasing their battery life.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

So part of my experiment explored how to teach the crew to do routine maintenance tasks, such as for the ATV brakes system. This is quite important, due to the safety considerations. I am aware that the ATVs all recently got a complete overhaul, which makes me grateful and confident in them, but I have one question. The brake fluid level indicators are barely readable, through a foggy window, so it’s hard to tell if there’s fluid in there at all. My question is then: were the brake systems checked, during the ATV checkout? Brake fluid is easy to replace, if needed.

Also, thank you all for supporting us throughout our mission! You’ve [CapComms and Mission Support] done a fantastic job at answering our questions and helping us with our concerns (especially reading my long explanations about that darned power system), and I’m very grateful!

Regards,

Melanie Grande, Crew Engineer, Crew 186

Sol Summary – January 12th

Crew 186 Sol 12 Summary Report 12JAN2018

Sol 12

Summary Title: A Bountiful Harvest Before Entry Interface

Author’s name: Max Fagin

Mission Status: Nearly complete…

Sol Activity Summary: After an animated debate over dimensional analysis and its applicability to breakfast foods (TLDR, we found a way for ‘The Pancake’ to be used as a non-SI unit of time, of distance, and of mass and energy; Manuscripts to be submitted to both Nature and Science upon our return) we geared up for the final EVA of our mission to the base of Skyline Rim. Once the EVA team had departed, Mark and Max began to harvest the microgreens from GreenHab. The EVA team returned with spectra and samples of the shales found in the region, along with a handful of shell fossils they had picked up on the way back through the Moons region. On Mars, such a discovery would be the scientific finding of the century. But here in the Moons region, they are so common as for it to be almost impossible to avoid stepping on them.

Once the sun had set, the team rotated through GreenHab and the Science dome collecting, weighing and cataloging our bountiful harvest of radish microgreens, green beans, cilantro, dill and lettuce. Finally, for our last meal in sim before our return to Earth, we subjected the food to its most critical experiment of all: Ingestion.

Look Ahead Plan: Tomorrow is handoff day. Looking forward to meeting the members of Crew 187 who are now en route. Safe travels!

Anomalies in work: None

Weather: 11C, Wind calm, Fair, Humidity 40%

Crew Physical Status: Healthy

EVA: Geological sampling of shales in the colluvial fan at the base of Skyline Ridge

Reports to be filed: Sol Summary, Operations Report, Science Report (Geology), Journalist Report, GreenHab Report, EVA Report, Mission Summary

Support Requested: None

Operations Report – January 11th

Crew 186 Operations Report 11JAN2018

SOL: 11

Name of person filing report: M. Grande

Non-nominal systems: None

Notes on non-nominal systems: None

Generator (hours run): 11.7h
Generator off, battery charging at 9:00am
Generator on at 6:40pm

Solar— SOC
@ 9:00am : 100%
@ 6:40pm : 56%

Diesel: 50%

Propane: 28%

Ethanol Free Gasoline (5 Gallon containers for ATV): 3.3 Gallons

Water (trailer): 10 Gallons

Water (static): 367 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 129380.8 Gallons

Toilet tank emptied: No

ATVs Used: None
Oil Added? No
ATV Fuel Used: 00 Gallons
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: Filled up the ATVs all the way, finally. Trying to track the fuel usage as well as the water usage for our sim, but it was difficult to see inside the fuel tank with the helmets. Now the fuel is right up to the neck of the tank, since we recently had the resupply, and we’re ready to roll for our long EVA tomorrow! One of them was making clanking sounds, and this will need to be inspected during our hand-off day.

Deimos rover used: No
Hours: 99.0
Beginning charge: 100%
Ending charge:
Currently charging: Yes

Sojourner rover used: Assigned to director only.
Hours: Director discretional hours
Beginning charge: 100%
Ending charge:
Currently charging: Yes

Spirit rover used: No
Hours: 10.5
Beginning charge: 100%
Ending charge:
Currently charging: Yes

Opportunity rover used: No
Hours: 5.5
Beginning charge: 100%
Ending charge:
Currently charging: No

Curiosity rover used: No
Hours: 3.1
Beginning charge: 100%
Ending charge:
Currently charging: No

HabCar used and why, where? No

General notes and comments: Today most of the crew didn’t go on EVA, so I used the time for my "Implications of VR for Crew Autonomy" experiment! I tested two crew members from my Pre-Mission Training Group and the Virtual Reality Training Group on tasks including ATV maintenance checks and geology sensing using a portable spectrometer. We used individual Engineering EVAs to make this happen. Lots of notes to write up tonight!

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support: Based on the unknown clanking sounds coming from one of the 350 ATVs, we won’t use that one tomorrow on EVA. Also, to confirm what I remember from training, the Diesel gauge needle ranges from "12:00 to 6:00, counterclockwise," right? The needle has stayed just over 50% ("9:00") for the past few days, even though we ran the generator for three days straight.

Regards,
Melanie Grande, Crew Engineer, Crew 186

Journalist Report – January 9th

[Sol 9]

Another slow day at the hab. The crew awoke to frigid temperatures and a shroud of blowing Martian dust – our first sandstorm. Unwilling to test our luck in the tumultuous conditions, we immediately cancelled the planned EVA and have postponed it to tomorrow. Despite the storm, however, life at the habitat remains quite pleasant. The rarefied Martian wind is too tenuous to threaten our immediate safety and instead fosters a sense of coziness here. The crew enjoyed a television show after breakfast and has spent the day working, reading, and debating various topics. For those who participated in yesterday’s chilly EVA the downtime was certainly welcome.

By the late afternoon the dust had begun to clear and a robotic supply rover which landed earlier this week was able to complete its slow trek to the hab. After a quick excursion to obtain the supplies the crew delightfully unpacked a brand new bread maker and put it to use right away. At the time of writing the team is eagerly awaiting the results.

Justin Mansell, MDRS Crew 186 Journalist

P.S. Photos attached. Photo of the day: 09Jan2018 Crew mental health questionable.jpg